OpenStudy (anonymous):

inverse of f(x)= 2x-5??

OpenStudy (anonymous):

y = 2x - 5 x = 2y - 5 x + 5 = 2y y = (x+5)/2 f^-(x) = (x+5)/2

OpenStudy (anonymous):

think of f(x) as y so y = 2x-5 now solve for x in terms of y: 2x = y+5 x = (y+5)/2